# Tight Bounds for Schrödinger Potential Estimation in Unpaired Data Translation

### v2

@@ G.1 General Implementation Details
Due to the strong restriction of the space in which the parameters are iterated to [1,1][-1,1] for both $b$ and $m$, it was possible to optimize them using Optuna [Akiba et al., 2019] and achieve good convergence of optimization. We used 50 trial optimization for optuna over the sliced $\mathbb{W}_{1}$ metric in the 25 Gaussian and I2I translation experiments and 100 trial optimization for optuna over the sliced $\mathbb{W}_{1}$ metric for Single Cell data.
@@
G.2 Metrics Used in the Experiments
@@
$v_{\theta}(y)=\sum_{k=1}^{K}\alpha_{k}\,\mathsf{p}(y;r_{k},\varepsilon S_{k}),\quad y\in\mathbb{R}^{d},$
   where, for any $k\in\{1,\dots,K\}$, $\mathsf{p}(\cdot\,;r_{k},\varepsilon S_{k})$ is the density of the Gaussian distribution ${\mathcal{N}}(r_{k},\varepsilon S_{k})$.
@@
9:Return the adjusted potential $v_{\theta}(y)=\sum\limits_{k=1}^{K}\alpha_{k}\,\mathsf{p}(y;r_{k},\varepsilon S_{k})$.
@@
    Provide quantitative, theoretically-grounded measures of distribution similarity
@@
    Are applicable to high-dimensional data
@@
    Can detect different failure modes of generative models (mode collapse, memorization, etc.)
@@
    Have been widely adopted in research papers and practical applications
$\text{Sliced}\penalty 10000\ \mathbb{W}_{1}(X,Y)=\frac{1}{N}\sum_{n=1}^{N}\mathbb{W}_{1}(\{\langle x_{i},\theta_{n}\rangle\}_{i=1}^{M},\{\langle y_{i},\theta_{n}\rangle\}_{i=1}^{M}),$
where $\{\theta_{n}\}_{n=1}^{N}$ are random projections uniformly distributed on $\mathbb{S}^{d-1}$.
with Gaussian kernel $k(x,y)=\exp(-\gamma\|x-y\|^{2})$.
where $X^{\prime}$ and $Y^{\prime}$ are independent copies.
$\text{Covered}_{k}=\mathbb{I}\left[\min_{x\in X}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})\leqslant\chi_{2,\alpha}^{2}\right],$
$\text{Covered Modes}=\sum_{k=1}^{K}\text{Covered}_{k}.$
@@
We also provide a description of the training algorithm for LightSB-OU - Algorithm 1. The pseudocode provided allows a reader to visually evaluate how our approach differs from regular LightSB.
G.3 Details of Evaluation on the Gaussian Mixture
@@
To find the metrics sliced $\mathbb{W}_{1}$, MMD, covered modes and energy distance, 5 iterations were used, including sampling and translation of 10000 random points from the standard normal distribution by trained models. They were used to find the mean values and standard deviations in the table and on the graphs.
@@
G.4 Details of Evaluation on the Unpaired Image-to-Image Translation
@@
G.5 Final Hyperparameter Values
@@
  • Parameters for experiment: $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=1e-3,\penalty 10000\ K=10,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=0.001,\penalty 10000\ m=-1.000.$
@@
  • Parameters for $T=1:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=1e-2,\penalty 10000\ K=100,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=0.0001,\penalty 10000\ m=-0.050,$
  • Parameters for $T=2:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=1e-2,\penalty 10000\ K=100,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=0.001,\penalty 10000\ m=-0.100,$
  • Parameters for $T=3:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=1e-2,\penalty 10000\ K=100,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=0.001,\penalty 10000\ m=-0.200.$
@@
It is worth noting that the process parameters $b,\penalty 10000\ m$ remained quite stable in all three experiments with a mixture of Gaussians with respect to changing the number of components $K$ for approximating the potential, which is in line with expectations. The process should facilitate transport independently of subsequent control, and therefore is likely to be unique and stable with respect to the hyperparameters in a well-posed problem.
Standard 25 Gaussians:
  • Parameters for $K=15:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=15,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.160,\penalty 10000\ m=-0.041,$
  • Parameters for $K=20:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=20,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.139,\penalty 10000\ m=-0.327,$
  • Parameters for $K=25:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=25,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.113,\penalty 10000\ m=-0.173,$
  • Parameters for $K=30:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=30,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.099,\penalty 10000\ m=-0.225.$
Irregular 25 Gaussians:
  • Parameters for $K=15:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=15,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.101,\penalty 10000\ m=-1.000,$
  • Parameters for $K=20:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=20,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.119,\penalty 10000\ m=-0.863,$
  • Parameters for $K=25:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=25,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.172,\penalty 10000\ m=-0.681,$
  • Parameters for $K=30:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=30,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.158,\penalty 10000\ m=-0.839.$
Anisotropic 25 Gaussians:
  • Parameters for $K=15:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=15,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.116,\penalty 10000\ m=-0.432,$
  • Parameters for $K=20:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=20,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.179,\penalty 10000\ m=-0.230,$
  • Parameters for $K=25:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=25,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.116,\penalty 10000\ m=-0.287,$
  • Parameters for $K=30:$ $\text{batch size}=128,\penalty 10000\ \varepsilon=0.1,\penalty 10000\ \text{lr}=2e-3,\penalty 10000\ K=30,\penalty 10000\ \text{diagonal}=\text{true},\penalty 10000\ b=-0.099,\penalty 10000\ m=-0.315.$

### v3

@@ F.1 General Implementation Details
Due to the strong restriction of the space in which the parameters are iterated to [1,1][-1,1] for both $b$ and $m$, it was possible to optimize them using Optuna (Akiba et al., 2019) and achieve good convergence of optimization. We used 50 trial optimization for optuna over the sliced $\mathbb{W}_{1}$ metric in the 25 Gaussian and I2I translation experiments and 100 trial optimization for optuna over the sliced $\mathbb{W}_{1}$ metric for Single Cell data.
@@
F.2 Metrics Used in the Experiments
@@
$v_{\theta}(y)=\sum_{k=1}^{K}\alpha_{k}\,\mathsf{p}(y;r_{k},\varepsilon\sigma_{T}^{2}S_{k}),\quad y\in\mathbb{R}^{d},$
   where, for any $k\in\{1,\dots,K\}$, $\mathsf{p}(\cdot\,;r_{k},\varepsilon\sigma_{T}^{2}S_{k})$ is the density of the Gaussian distribution ${\mathcal{N}}(r_{k},\varepsilon\sigma_{T}^{2}S_{k})$.
@@
9:Return the adjusted potential $v_{\theta}(y)=\sum\limits_{k=1}^{K}\alpha_{k}\,\mathsf{p}(y;r_{k},\varepsilon\sigma_{T}^{2}S_{k})$.
@@
  1. (a)
    provide quantitative, theoretically-grounded measures of distribution similarity;
  2. (b)
    are applicable to high-dimensional data;
  3. (c)
    can detect different failure modes of generative models (mode collapse, memorization, etc.);
  4. (d)
    have been widely adopted in research papers and practical applications.
Metrics used in the empirical analysis of LightSB-OU:
@@
    Sliced Wasserstein Distance:
    $\text{Sliced}~\mathbb{W}_{1}(X,Y)=\frac{1}{N}\sum_{n=1}^{N}\mathbb{W}_{1}\left(\{\langle x_{i},\theta_{n}\rangle\}_{i=1}^{M},\{\langle y_{i},\theta_{n}\rangle\}_{i=1}^{M}\right),$
    where $\{\theta_{n}:1\leqslant n\leqslant N\}$ are i.i.d. random vectors drawn from the uniform distributed on the unit sphere $\mathbb{S}^{d-1}$.
@@
    Maximum Mean Discrepancy:
    with Gaussian kernel $k(x,y)=\exp\{-\gamma\|x-y\|^{2}\}$.
@@
    Energy Distance:
    where $X^{\prime}$ and $Y^{\prime}$ are independent copies of $X$ and $Y$, respectively.
@@
    Number of Covered Modes (for Gaussian Mixture Experiment):
    $\text{Covered Modes}=\sum_{k=1}^{K}\text{Covered}_{k},$
    where
    $\text{Covered}_{k}=\mathbb{I}\left[\min_{x\in X}(x-\mu_{k})^{\top}\Sigma_{k}^{-1}(x-\mu_{k})\leqslant\chi_{2,\alpha}^{2}\right].$
@@
We also provide a description of the training algorithm for LightSB-OU (Algorithm 1). The pseudocode provided allows a reader to visually evaluate how our approach differs from regular LightSB.
F.3 Details of Evaluation on the Gaussian Mixture
@@
To find the metrics sliced $\mathbb{W}_{1}$, MMD, covered modes and energy distance, 5 iterations were used, including sampling and translation of 10000 random points from the standard normal distribution by trained models. They were used to find the mean values and standard deviations in the table and on the graphs. During the inference, we have used the fact that the conditional density of $X_{T}^{\varphi_{\theta}}$ given $X_{0}^{\varphi_{\theta}}=x$ has a form (see (20))
$\pi^{\varphi_{\theta}}\left(y\,\big|\,x\right)=\frac{\pi^{\varphi_{\theta}}(x,y)}{\rho_{0}(x)}=e^{y^{\top}m_{1}(x)/(\varepsilon\sigma_{1}^{2})}v_{\theta}(y)\big/\penalty 50c_{\theta}(x),$
where $v_{\theta}(y)$ and $c_{\theta}(x)$ are given by (19) and (21), respectively. When $v_{\theta}(y)$ is a Gaussian mixture, it is straightforward to check that the conditional distribution is a Gaussian mixture as well. More specifically, if $v_{\theta}$ is defined by (19), then
$\pi^{\varphi_{\theta}}\left(y\,\big|\,x\right)=\sum\limits_{k=1}^{K}\widetilde{\alpha}_{k}(x)\mathsf{p}(y;r_{k}(x),\varepsilon\sigma_{1}^{2}S_{k}),$
where $\widetilde{\alpha}_{k}(x)$ is given by (22), $\mathsf{p}(y;r_{k}(x),\varepsilon\sigma_{1}^{2}S_{k})$ stands for the density of ${\mathcal{N}}\big(r_{k}(x),\varepsilon\sigma_{1}^{2}S_{k}\big)$ and $r_{k}(x)=r_{k}+S_{k}m_{1}(x)$ for all $k\in\{1,\dots,K\}$.
@@
Figure 5: The translation result from a standard normal distribution using LightSB (left) and our modified approach (right). Top: uniform grid and standard covariance matrices. Middle: random location on the grid and standard covariance matrices. Bottom: uniform grid and anisotropic random covariance matrix. Green: samples from the target Gaussian mixture; orange: LightSB samples; blue: LightSB-OU samples.
F.4 Details of Evaluation on the Single Cell Data
We analyzed cell differentiation across five stages from day 0 to 27 (i.e. $t_{0}:$ day 0 to 3, $t_{1}:$ day 6 to 9, $t_{2}:$ day 12 to 15, $t_{3}:$ day 18 to 21, $t_{4}:$ day 24 to 27), processing the scRNA-seq data through quality filters and PCA to create distinct feature vectors. In this experiment the goal was to predict the cell distribution at an intermediate time point ($t_{i}$) by ”transporting” data from the surrounding intervals ($t_{i-1}$ and $t_{i+1}$), where $i\in[1,2,3].$ To evaluate the model’s accuracy, we calculated the Wasserstein-1 ($\mathbb{W}_{1}$) distance to see how closely our predictions matched the actual observed data.
F.5 Details of Evaluation on the Unpaired Image-to-Image Translation
@@
F.6 Final Hyperparameter Values
@@
Empirical KL Divergence Experiment:
  • Parameters for experiment: $\text{batch size}=128,~\varepsilon=9.0,~\text{lr}=0.002,~K=1,~\text{diagonal}=\text{false},~b=2.0,~m=0.0.$
@@
  • Parameters for experiment: $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.001,~K=10,~\text{diagonal}=\text{true},~b=0.02,~m=0.0.$
@@
  • Parameters for $i=1:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.01,~K=100,~\text{diagonal}=\text{true},~b=-0.2,~m=4.0,$
  • Parameters for $i=2:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.01,~K=100,~\text{diagonal}=\text{true},~b=-0.2,~m=4.0,$
  • Parameters for $i=3:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.01,~K=100,~\text{diagonal}=\text{true},~b=0.2,~m=-1.0.$
@@
It is worth noting that the process parameters b,mb,~m remained quite stable in all three experiments with a mixture of Gaussians with respect to changing the number of components $K$ for approximating the potential, which is in line with expectations. The process should facilitate transport independently of subsequent control, and therefore is likely to be unique and stable with respect to the hyperparameters in a well-posed problem.
Standard Gaussian mixture with 25 components:
  • Parameters for $K=15:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=15,~\text{diagonal}=\text{true},~b=-0.125,~m=0.901,$
  • Parameters for $K=20:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=20,~\text{diagonal}=\text{true},~b=-0.125,~m=0.901,$
  • Parameters for $K=25:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=25,~\text{diagonal}=\text{true},~b=0.232,~m=0.197,$
  • Parameters for $K=30:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=30,~\text{diagonal}=\text{true},~b=0.101,~m=0.416.$
Irregular Gaussian mixture with 25 components:
  • Parameters for $K=15:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=15,~\text{diagonal}=\text{true},~b=-0.125,~m=0.901,$
  • Parameters for $K=20:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=20,~\text{diagonal}=\text{true},~b=0.332,~m=-0.575,$
  • Parameters for $K=25:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=25,~\text{diagonal}=\text{true},~b=0.332,~m=-0.575,$
  • Parameters for $K=30:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=30,~\text{diagonal}=\text{true},~b=0.332,~m=-0.575.$
Anisotropic Gaussian mixture with 25 components:
  • Parameters for $K=15:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=15,~\text{diagonal}=\text{true},~b=-0.125,~m=0.901,$
  • Parameters for $K=20:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=20,~\text{diagonal}=\text{true},~b=-0.125,~m=0.901,$
  • Parameters for $K=25:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=25,~\text{diagonal}=\text{true},~b=0.232,~m=0.197,$
  • Parameters for $K=30:$ $\text{batch size}=128,~\varepsilon=0.1,~\text{lr}=0.002,~K=30,~\text{diagonal}=\text{true},~b=0.101,~m=0.416.$